OpenStudy (anonymous):

Factor completely: 2x2 + 9x + 4 please explain it :(

Parth (parthkohli):

Find two numbers a,b such that: a * b = 2 * 4 = 8 and a + b = 9

Parth (parthkohli):

@moosic?